OpenStudy (anonymous):

Score: Match the following items by evaluating the expression for x = -2. x^(-2)

OpenStudy (anonymous):

@jim_thompson5910 will u help me

jimthompson5910 (jim_thompson5910):

x^(-2) = 1/( x^2 ) now plug in x = -2 and evaluate

OpenStudy (anonymous):

so I got -1/2

jimthompson5910 (jim_thompson5910):

(-2)^2 = ???

jimthompson5910 (jim_thompson5910):

what do you get when you square -2

OpenStudy (anonymous):

4

jimthompson5910 (jim_thompson5910):

1/(x^2) = 1/( (-2)^2 ) 1/(x^2) = 1/4

jimthompson5910 (jim_thompson5910):

so 1/(x^2) = 1/4 when x = -2